OpenStudy (kpop4life123):

If x3 = 216, what is the value of x? (5 points) 4 3 9 6

OpenStudy (whpalmer4):

\[x^3=216\]is the same as \[x*x*x=216\] which of those numbers substituted for \(x\) makes a true statement? Hint:\[3*3*3 = 27\]so it isn't \(3\)

OpenStudy (kpop4life123):

6!